OpenStudy (anonymous):

need help with this question 2.5x+y=-2 3x+2y=0

OpenStudy (anonymous):

You can solve this system of equation using the elimination or substitution method.

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

Do you know either method?

OpenStudy (anonymous):

i have but i keep getting the wrong answer i dont know what im doing wrong

OpenStudy (anonymous):

can you walk me through it?

OpenStudy (anonymous):

Alright, multiply top eqution by -2 -2(2.5x+y=-2)..... -5x-2y=2 leave the second eqution the same 3x+2y=0

OpenStudy (anonymous):

that's not the way i was tought to do it ...

OpenStudy (anonymous):

Show me then.

OpenStudy (anonymous):

-5x-2y=4 3x+2y=0 eliminate the y variable -2x=4 x=-2 Now plug this back into any equation.

OpenStudy (anonymous):

3(-2)+2y=0 -6+2y=0 2y=6 y=3 therefore, x=-2, y=3

OpenStudy (anonymous):

Does that help?